OpenStudy (tomektomeknyc):

how to solve this: sin theta cot ^2 theta sec ^2 theta The answers: a) sec theta b) csc theta c) tan theta sec theta d) cos theta

OpenStudy (kc_kennylau):

Do you mean:\[\sin\theta\cot^2\theta\sec^2\theta\]

OpenStudy (kc_kennylau):

Hint: cot=cos/sin and sec=1/cos :)

OpenStudy (anonymous):

.